OpenStudy (anonymous):

Since the Celsius is less then or equal to the Fahrenheit Using F= (9/5)C + 32 , we now have F = (9/5)-111.4 +32 F = (9/5)-111.4 +32 F = 9/5 *5/9 -111.4 + 32 F= -79.4 Check: -79.4= 9/5 -111.4 +32 -79.4=-79.4 True is this correct?

OpenStudy (sriram):

first of all i dont think celsius is always less than equal to fahrenheit and in the equation are u trying to substitute 111.4 for C?

OpenStudy (anonymous):

yes that what the problem stated. In order for a chemical reaction to remain stable, its Celsius temperature must be no more than 111.4 degrees Celsius. Find the Fahrenheit temperatures at which the reaction will remain stable.

OpenStudy (anonymous):

so this becomes undefined? Or can it be solved?

OpenStudy (sriram):

acc. to ur equation on calculating F we get 232.52

OpenStudy (sriram):

celcius and fahrenheit are equal at -40

OpenStudy (anonymous):

How did u get that and What did I do wrong in my steps?

OpenStudy (sriram):

use a calculator 9/5 *111.4=200.52 200.52+32=232.52

OpenStudy (anonymous):

ok so i was to multiply not change the sign to negative. thanks a lot for the tip. I can solve it correctly now. thanks medal given

OpenStudy (anonymous):

F = C * 9/5 + 32 F -32 = C * 9/5 + 32 -32 - Adding the additive inverse of a number to both sides of the equation to move it to the other side of it F -32 = C * 9/5 - Addition ( F -32) * 5/9 = C * 9/5 * 5/9 - Multiplying the multiplicative inverse of a number to both sides of the equation to move it to the other side of it 5/9( F-32) = C - Multiplication C = 5/9( F- 32) - Rearranged the equation around 111.4 = 5/9 ( 232.52-32) - Substitution of C and F with 111.4 and 232.52 111.4 = 5/9( 200.52) - Subtraction 111.4 = 1,002.60 / 9 - Multiplication 111.4 = 111.4 - Division It checks and equals therefore the chemical reaction will remain stable, as long as it temperature in Farenheit is less than or equal to 232.52 degrees which is 111.4 degrees Celsius